# The NTU Partitioned Matching Game for International Kidney Exchange Programs

### v1

@@ Abstract
Motivated by the real-world problem of international kidney exchange (IKEP), [Biró et al., Generalized Matching Games for International Kidney Exchange, 2019] introduced a generalized transferable utility matching game featuring a partition of the vertex set of a graph into players, and analyzed its complexity. We explore the non-transferable utility (NTU) variant of the game, where the utility of players is given by the number of their matched vertices. The NTU version is arguably a more natural model of the international kidney exchange program, as the utility of a participating country mostly depends on how many of its patients receive a kidney, which is non-transferable by nature. We study the core of this game, which suitably captures the notion of stability of an IKEP, as it precludes incentives to deviate from the proposed solution for any possible coalition of the players.
We prove computational complexity results about the weak and strong cores under various assumptions on the players. In particular, we show that if every player has two vertices, which can be considered as an NTU matching game with couples, then the weak core is always non-empty, and the existence of a strong core solution can be decided in polynomial time. In contrast, it is NP-hard to decide whether the strong core is empty when each player has three vertices. We also show that if the number of players is constant, then the non-emptiness of the weak and strong cores is polynomial-time decidable.
@@ 1 Introduction
In kidney exchange programs (KEP), patients may swap their incompatible donors, leading to cycles of kidney transplants [5]. In order to increase the availability of compatible donors, it is desirable for countries to merge their national patient-donor pools, leading to international KEPs (IKEPs) [7, 16]. A joint pool of multiple countries helps more patients receive kidneys and can also improve the average compatibility of transplants. However, to keep the cooperation mutually beneficial for all participating countries, it is important to guarantee some form of stability and fairness, at least in the long run. Hence, it is natural to study IKEPs in the context of game theory, which provides several theoretical tools for the analysis of various forms of stability and fairness.
In the framework of cooperative game theory, an IKEP is considered as a game where the players are the countries, and each country’s main objective is to match as many of its patients to some compatible donors as possible (in a more general model, the utility of transplantations can be characterized by edge weights, but in this paper we only consider the unweighted version). Due to difficulties and dangers involved in organizing large cycles of transplantations, it is usual to allow pairwise exchanges only. Possible pairwise exchanges are modeled by a graph $G=(V,E)$, where the vertices represent the patient-donor pairs, and an edge $uv$ represents a mutually compatible exchange. The aim of the IKEP is then to determine a matching in this graph, and each country’s main goal is for this matching to cover as many of its patients (i.e. the vertices corresponding to the patient-donor pairs of the country) as possible.
Previous literature [11, 12, 14] suggested to maintain a level of fairness using a scoring system in a multi-round mechanism. In [12], the concept of fair allocations in transferable utility (TU) games is used in each round of the IKEP to get a target allocation. This approach uses the TU version of the cooperative partitioned matching game. A comprehensive analysis of the computational aspects can be found in [3].
In the present paper, we study the computational complexity of the non-transferable utility (NTU) version of partitioned matching games. We argue that the NTU version models IKEPs more naturally, as the utilities of the countries in these programs are best described by the number of their patients that receive a kidney, which is non-transferable by nature.
In the NTU setting, we are interested in solutions where no coalition of countries has an incentive to deviate from the proposed solution in order to match more of their patients. In particular, we consider the well-known concepts of strong core and weak core in NTU games. That is, we are looking for solutions where no coalition of the players (i.e. the countries) could create a solution among themselves where each of them improves (in case of the weak core), or some of them improves and none of them has fewer patients matched (in case of the strong core).
Players in a partitioned matching game may represent entities other than countries. For example, in the context of national kidney exchange programs (KEPs), hospitals can be considered as the players of the game, which gives rise to instances where each player (hospital) has only a few vertices of the graph (patients). Thus, it is meaningful to examine the computational complexity of the problem when each player has a bounded number of vertices.
In particular, it is interesting to consider the special case where each player has only two vertices, i.e., the players correspond to couples. This is a new kind of matching problem with couples, which can be a useful modeling tool outside the realm of kidney exchange. For example, take a situation where some couples get together to play a two-person sport like squash and have to form pairs that play against each other. Each person has a list of possible opponents that they are willing to play against (which may or may not include their partner). If the graph of feasible pairs has a perfect matching, then everyone can play. Otherwise, the best option for a couple is when both can play, and the worst option is when neither of them can. We want to find a pairing where no subset of couples can organize the matches better among themselves. This gives rise to an NTU partitioned matching game with partition classes of size 2. We call this special case the NTU matching game with couples. As another motivating example, we may consider a summer research exchange program where couples are student-supervisor pairs, and we want to match each student to another supervisor whose topic is interesting to the student. The best option for an original student-supervisor pair is when both are matched in the program, and the worst is when neither are matched. Note that the graph describing the possible matches is bipartite in this instance.
@@
The different contexts mentioned above justify the study of the computational complexity both for instances where the size of each partition class is bounded, and for instances where the number of partition classes (i.e., players) is bounded. Our main results are as follows:
@@
  • For the NTU matching game with couples, we show that the weak core is always non-empty and a weak core solution can be found in polynomial time. As the technically most involved result of the paper, we construct a polynomial-time algorithm for deciding if the strong core is empty, and for finding a strong core solution, if it exists. The proof uses a complicated analysis of the structure of paths and cycles that are alternating between edges of the graph and “edges” corresponding to the couples. An interesting consequence of our result is that the vertex sets of matchings in the strong core can be described as the common bases of two matroids (it is well known that vertex sets of maximum matchings are the bases of the so-called matching matroid; matchings in the strong core are additionally bases of a generalized partition matroid defined by the structure of alternating paths and cycles).
@@
  • We show that if the number of players is bounded by a constant, then the problems of checking membership in the weak and strong cores, and of finding elements in the weak and strong cores are all polynomial-time solvable. These results follow fairly easily using tools from matroid theory; here we use common bases of a matching matroid and a generalized partition matroid in a different context.
@@
  • In contrast to the above results, we show that problems where the size of each partition class is bounded by a constant are hard even for fairly small constants. In particular, deciding whether a matching is in the weak or strong core is coNP-complete even when each partition class has size at most 3.
@@
Our results provide useful new tools for designing mechanisms for IKEPs. In particular, as most IKEPs consist of only a few countries, our results imply that we can efficiently find strong and weak core solutions for the NTU version of the partitioned matching game in those cases. Such solutions, if they exist, give desirable outcomes for exchange programs, as strategic incentives to deviate from the program can be avoided. Also, as membership of a matching in the strong or weak core depends only on the number of vertices it covers from each player, a nonempty core always contains a maximum size matching (this follows from the fact that any set of vertices coverable by a matching is also coverable by a maximum size matching). Therefore, our concepts are compatible with the most important restriction used in KEPs, namely that the number of transplants should be maximized.
@@
In addition to the applicability of our results to kidney exchange problems, our algorithm for the problem with couples offers a new class of non-bipartite matching games where stable solutions can be efficiently found even in the presence of couples having joint preferences.
@@
In Section 2, we show how membership in the weak and strong cores in the NTU matching game with couples can be characterized using alternating paths. We then prove that the weak core is always non-empty in this case. The strong core can be empty, but we are able to give a structural characterization and a polynomial-time algorithm for finding a strong core solution if it exists.
@@
In Section 3, we show that all problems studied are polynomial-time solvable if the number of players is constant, even if the sizes of the partition classes can be arbitrarily large. Section 4 presents hardess results for the case where the partition classes have bounded size. We show that all problems are hard for some small constant bound, and for most problems we can prove coNP-hardness even for partition classes of size at most 3.
Related work
In [1, 2], Ashlagi and Roth considered individual rationality in the multi-hospital kidney exchange problem. In terms of NTU games, individual rationality means that there is no single blocking player. They also studied the problem of incentive-compatibility, i.e., whether hospitals are motivated to underreport their patient-donor pairs. Carvalho et al. [8] studied the properties of Nash-equilibria in the two-hospital case. Gourvès, Monnot and Pascual [10] considered a weighted bipartite graph model, where the profit of an edge in the matching is divided between the two endpoints according to a fixed ratio. They proved that it is NP-hard to decide if there is an individually rational maximum weight matching.
As mentioned in the introduction, credit-based multi-round mechanisms for IKEPs were studied in [11, 12, 14]. Solution concepts of cooperative game theory were used in [12] to find target allocations in each round. The mechanism then computes a maximum size matching that approximates this target allocation for each country. At the end of the round, the scores are updated according to whether the number of matched patients of a country is above or below the target. Computational experiments on large-scale instances were performed in [4], while [3] studied the theoretical computational complexity of the problems involved. In particular, they examined problems similar to ours in the transferable utility setting, and showed that checking core non-emptiness in TU partitioned matching games is polynomial-time solvable for partition classes of size at most 2, but co-NP-hard for partition classes of size at most 3.
We note that a previous version of the present paper, including only a subset of the results, appeared as a technical report [13]. Independently of our work, the Master’s thesis of Collette [9] also studied the NTU version of the partitioned matching game, in particular the weak core. It offered a cutting plane approach to compute a weak core solution, as well as a comparison with the rejection-proof kidney exchange mechanism of Blom et al. [6], and computational experiments.
Preliminaries

### v2

@@ Abstract
Motivated by the real-world problem of international kidney exchange (IKEP), recent literature introduced a generalized transferable utility matching game featuring a partition of the vertex set of a graph into players, and analyzed its complexity. We explore the non-transferable utility (NTU) variant of the game, where the utility of players is given by the number of their matched vertices. Our motivation for studying this problem is twofold. First, the NTU version is arguably a more natural model of the international kidney exchange program, as the utility of a participating country mostly depends on how many of its patients receive a kidney, which is non-transferable by nature. Second, the special case where each player has two vertices, which we call the NTU matching game with couples, is interesting in its own right and has intriguing structural properties.
We study the core of the NTU game, which suitably captures the notion of stability of an IKEP, as it precludes incentives to deviate from the proposed solution for any possible coalition of the players. We prove computational complexity results about the weak and strong cores under various assumptions on the players. In particular, we show that if every player has two vertices, then the weak core is always nonempty, and the existence of a strong core solution can be decided in polynomial time. Moreover, one can efficiently optimize on the strong core. In contrast, it is NP-hard to decide whether the strong core is empty when each player has three vertices. We also show that if the number of players is constant, then the non-emptiness of the weak and strong cores is polynomial-time decidable, and we can find a minimum-cost core solution in polynomial time.
Keywords— graph algorithms, matching, cooperative game, kidney exchange
@@ 1 Introduction
We study the NTU Partitioned Matching Game, which is a cooperative game defined as follows. Let $G=(V;E)$ be a graph, and let $V_{1},V_{2},\dots,V_{m}$ be a partition of $V$ into $m$ subsets, where $m$ is the number of players. Each class of this partition corresponds to a player. The utility of a subset $\{i_{1},\dots,i_{k}\}$ of players is defined as the maximum size of a matching in the subgraph of $G$ induced by the vertex set $V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}$. The transferable utility (TU) version of this game was introduced by Kern et al. [31]. The non-transferable utility (NTU) version, where the utility of a matching $M$ for player $i$ is the number of vertices of $V_{i}$ covered by $M$, has been considered in a technical report by a subset of the present authors [32], as well as by Collette [23].
In this paper, we focus on computational complexity problems related to the weak and strong cores of the NTU game. Our motivation for studying the core is twofold. First, International Kidney Exchange Programs (IKEPs) can be naturally modeled as a partitioned matching game, and considering the NTU core is an appealing way to strengthen the concept of individual rationality, which takes into account the possibility that some countries might have the incentive to abandon the program together and arrange the exchanges among themselves. The second motivation comes from theoretical considerations: the computational complexity of the core has particularly interesting properties in the case where every partition class has size at most 2, which we call the NTU matching game with couples. In particular, we can prove that the non-emptiness of the strong core can be decided in polynomial time. The proof is quite involved and requires complex arguments from matching theory. In the following, we describe these two motivations in detail.
1.1 Motivation: international kidney exchange programs
Kidney exchange programs (KEPs) enable patients with incompatible live donors to receive transplants by swapping donors with other incompatible pairs. These exchanges, typically organized into 2-way or 3-way cycles, have facilitated thousands of life-saving transplants worldwide [14]. To improve access and efficiency, some national programs have begun to collaborate through international kidney exchange programs (IKEPs), pooling patient-donor pairs across borders [20, 49]. While international coordination substantially expands the set of potential matches and can increase both the number and quality of transplants, it also introduces new challenges — most notably, how to incentivize countries (or hospitals, for example in the US setting) to actively participate in these programs.
These incentive issues are not merely theoretical. In the US, the kidney exchange market facilitates hundreds of transplants annually, yet remains fragmented and inefficient [1]. Hospitals often withhold easy-to-match pairs, participating only with hard-to-match patients, which undermines the overall efficiency of the system. A similar pattern appears in Europe, where IKEPs are often limited to hard-to-match pairs that were not matched during national kidney exchange programs [14], greatly reducing their effectiveness [25]. Institutional fragmentation and misaligned incentives therefore remain a central obstacle to improving outcomes in both national and international kidney exchange settings.
To avoid serious challenges that arise when trying to ensure strategy-proofness (in particular, the impossibility of achieving both strategy-proofness and maximality or near-maximality of the number of transplants [4, 5, 40]), in this paper we adopt the framework of cooperative game theory, modeling IKEPs as partitioned matching games where players – typically countries or hospitals– seek to maximize the number of their own patients who receive a transplant, similarly as in [31]. We focus on pairwise exchanges (2-cycles), and assume unweighted utility, where each matched patient contributes equally to the country’s objective. Due to medical and logistical constraints, allowing only pairwise exchanges is adopted by many countries, such as Denmark, Norway, Sweden and France [14, 16], where the former three also share cooperation internationally. The underlying matching problem is represented by an undirected graph $G=(V,E)$, where vertices denote patient-donor pairs and edges indicate mutually compatible exchanges.
As opposed to Kern et al. [31], in this paper we focus on the non-transferable utility (NTU) setting, which is a natural fit for kidney exchange: the utility (number of patients matched) cannot be transferred among players. This is also in line with how the literature for the US market models the utilities of the participants [3, 6, 5, 40, 41]. We investigate the existence and computational complexity of core solutions in NTU partitioned matching games, particularly exploring two important stability notions: the strong core, where no coalition can weakly block the solution (meaning only one participant has to improve), and the weak core, where no coalition can strongly block (i.e. all participants must improve) it.
Beyond the international setting, our framework also applies to national KEPs like the US setting, where the players may represent hospitals rather than countries [46]. In such cases, each institution typically controls only a small number of patient-donor pairs. This motivates the study of partitioned matching games where each player controls a bounded number of vertices.
1.2 Motivation: NTU matching game with couples
The NTU Partitioned Matching Game has many remarkable properties also from the point of view of algorithmic game theory and computational complexity. It turns out that the most interesting case from this point of view is when each player controls two vertices of the graph, which we call the NTU matching game with couples (we remark that the case where each partition class has size at most two can be easily reduced to the case where each partition class has size exactly two, by introducing dummy vertices). Although this case is not very relevant to the kidney exchange problem, we believe that it is interesting on its own right, and its analysis requires sophisticated arguments based on matching theory. To have an intuitive understanding of this game, we can imagine some couples who come together to play some sport, for example squash, with two-player matches. Each individual has binary preferences, i.e. there are people they like to play with, and there are those they prefer not to play with (we assume that these preferences are mutual, and of course it is possible that a couple prefers not to play with each other). The task is to organize the matches in a way that is acceptable to all couples, where the best option for a couple is that both of them play with a preferred partner, and the second best option is that one of them plays with a preferred partner. In this setting, the core is a natural and intuitive solution concept, because it means that no subset of couples has the incentive to leave the group and organize their matches among themselves.
As we will see, the weak core is always non-empty in this case, but the non-emptiness of the strong core is an intriguing question that is intimately related to the structure of paths that alternate between edges of the graph and “edges” corresponding to couples. Furthermore, although the previous statement suggests that the structure of the weak core is simpler than that of the strong core, this is in fact not the case: it turns out that optimizing a linear function on the weak core is NP-hard, while the same problem on the strong core is solvable in polynomial time.
Our results on this problem fit into a long sequence of results on matching problems with couples, where the introduction of joint preferences of couples leads to intriguing complexity questions [38, 2, 33, 35, 36, 12, 37, 24].
1.3 Our results
@@
The motivations mentioned above justify the study of the computational complexity both for instances where the size of each partition class is bounded, and for instances where the number of partition classes (i.e., players) is bounded. Our main results are as follows:
@@
  • For the NTU matching game with couples, we show that the weak core is always nonempty and a weak core solution can be found in polynomial time. As the technically most involved result of the paper, we construct a polynomial-time algorithm for deciding if the strong core is empty, and for finding a strong core solution, if it exists. The proof uses an analysis of the structure of paths and cycles that are alternating between edges of the graph and “edges” corresponding to the couples. Our results also imply that we can efficiently optimize on the strong core, i.e., for any given edge costs, we can find a minimum cost matching in the strong core in polynomial time. This is remarkable since optimizing on the weak core turns out to be an NP-hard problem.
@@
  • We show that if the number of players is bounded by a constant, then the problems of checking membership in the weak and strong cores, and of finding minimum cost elements in the weak and strong cores are all polynomial-time solvable.
@@
  • In contrast to the above results, we show that problems where the size of each partition class is bounded by a constant are hard even for fairly small constants. In particular, deciding whether a matching is in the weak or strong core is coNP-complete even when each partition class has size at most 3. Similarly, deciding the weak or strong core’s emptiness also becomes NP-hard, even for small sizes.
@@
Our results provide useful new tools for designing mechanisms for IKEPs. In particular, as most IKEPs consist of only a few countries, our results imply that we can efficiently find strong and weak core solutions for the NTU version of the partitioned matching game in those cases. Such solutions, if they exist, give desirable outcomes for exchange programs, as incentives to quit the program can be avoided. Also, as membership of a matching in the strong or weak core depends only on the number of vertices it covers from each player, a nonempty core always contains a maximum size matching (this follows from the fact that any set of vertices coverable by a matching is also coverable by a maximum size matching). Therefore, our concepts are compatible with the most important restriction used in KEPs, namely that the number of transplants should be maximized. Furthermore, as a simulation study of Colette [23] – using the state-of-the-art KEP generator by Saidman et al. [43] – recently showed, the weak core of realistic international kidney exchange markets always admits a weak core exchange. Hence, even though existence is not guaranteed in theory, instances with an empty weak core arise only in special constructions, but not in real-world markets. Furthermore, this also highlights that even though most arising problems are NP-hard in theory, it is feasible to compute core solutions in practice with the help of integer programming techniques.
@@
In addition to the applicability of our results to kidney exchange problems, our algorithm for the problem with couples offers a new class of non-bipartite matching games where core solutions can be efficiently found even in the presence of couples having joint preferences.
@@
In Section 2, we show how membership in the weak and strong cores in the NTU matching game with couples can be characterized using alternating paths. We then prove that the weak core is always non-empty in this case; however, optimizing on the weak core is NP-hard. The strong core can be empty, but we are able to give a structural characterization and a polynomial-time algorithm for finding a minimum cost strong core solution if it exists.
@@
In Section 3, we show that all problems studied are polynomial-time solvable if the number of players is constant, even if the sizes of the partition classes can be arbitrarily large. Section 4 presents hardness results for the case where the partition classes have bounded size. We show that all problems are hard for some small constant bound, and for most problems we can prove coNP-hardness even for partition classes of size at most 3.
1.4 Related work
Transferable Utility (TU) Matching Games.
Matching games with transferable utilities are known as assignment games. A foundational result by Shapley and Shubik [45] shows that the core of any assignment game is non-empty. This result was extended to many-to-many settings by Sotomayor [47], who proved core non-emptiness for any capacity function $b$.
Computational aspects of core-related problems have also been studied. Biró et al. [15] showed that deciding whether an allocation lies in the core is coNP-complete even for uniform 3-assignment games. For $b\leq 2$, however, they proved that both core verification and finding a core solution are solvable in polynomial time.
The partitioned matching game was introduced by Kern et al. [31] to model international kidney exchange programs (IKEPs), where players control subsets of vertices rather than individual ones. Later works [10, 9] examined the computational complexity of these games and showed that the core can be empty. They also established NP-hardness and coNP-hardness for core non-emptiness and verification, even when players control at most three vertices. A close relationship was established with many-to-many assignment games via two-way reductions.
A variant of these games involving unbounded exchanges, called partitioned permutation games, was studied by Benedek et al. [7]. They demonstrated that such games always admit a core solution, which can be found in polynomial time. Nevertheless, verifying whether a given solution belongs to the core remains coNP-complete.
Non-Transferable Utility (NTU) Matching Games.
In the NTU setting, Scarf’s classical result [44] establishes the non-emptiness of the (weak) core in a broad class of games, called balanced games. Kaneko [30] applied Scarf’s Lemma to the NTU version of the Shapley-Shubik assignment game and showed that the weak core is always non-empty, even in general settings. Biró and Fleiner [13] further used Scarf’s Lemma to establish the existence of fractional weak core solutions in hypergraph matching games.
Stability is a concept closely related to the core. In many-to-one matching markets, Roth [42] showed that under strict preferences, the strong core coincides with stable matchings. Gale and Shapley [26] proved the existence of stable matchings in bipartite graphs, while Irving [29] gave a polynomial algorithm for non-bipartite settings with strict preferences, where existence is no longer guaranteed.
However, in many-to-many markets, the connection between core and stability weakens. Blair [17] and Sotomayor [48] demonstrated that the strong core may be empty and does not coincide with stable outcomes. Recently, Biró and Csáji [11] showed that deciding strong core non-emptiness is NP-hard even in restricted many-to-many markets, and that verifying core membership is coNP-hard.
The NTU version of the partitioned matching game was first explored in our earlier technical report [32], which established basic complexity results. Independently, Collette’s Master’s thesis [23] also studied the weak core in NTU partitioned matching games. He proved that the weak core is always non-empty for two players, and provided a no-instance for the core with 3 players - with 9 vertices each, compared to our no-instance with 7 vertices per player. Then, he proposed a (not polynomial) cutting-plane-based algorithm to compute weak core outcomes. Using the Saidman generator [43], he showed empirically that weak core matchings always existed in simulations, highlighting the practical viability of this concept. He also compared weak core outcomes with those of rejection-proof mechanisms proposed by Blom et al. [18].
Incentives in kidney exchange programs.
The literature on incentive issues in kidney exchange programs can be broadly classified into three main streams, depending on whether they adopt a cooperative, non-cooperative, or mechanism-design perspective. We summarize these below and highlight their relationship to our approach.
Cooperative approaches. In the European setting, a common approach is to treat countries as cooperating players and to focus on fairness over time across multiple exchange rounds [28, 31, 34]. In these systems, each country is assigned a “target” number of transplants in each round based on a fairness criterion. Because the actual matches in a round may deviate from these targets due to medical and logistical constraints, discrepancies are tracked as credits (positive or negative) and carried forward to future rounds. Over time, this mechanism balances outcomes across rounds. Kern et al. [31] proposed using solution concepts from transferable utility (TU) cooperative game theory to determine these target allocations, such as the benefit value, contribution value, Shapley value, or Banzhaf value. A key advantage of this approach is that it is compatible with maximum-size matching: in each round, the actual matching is chosen to maximize the total number of transplants, and fairness is addressed through the credit adjustments across rounds.
Non-cooperative approaches. In the US setting, hospitals are typically modeled as self-interested players in a non-cooperative game. Much of the literature has focused on strategy-proofness and individual rationality (IR), where IR in NTU terms corresponds to the absence of a single blocking player [6, 4, 5]. A major challenge in this setting is that strong incentive guarantees come at a cost to efficiency. Ashlagi and Roth [5] proved that no mechanism can be both strategy-proof and individually rational while guaranteeing more than half of the maximum possible number of transplants, even when only 2-way cycles are allowed. Ashlagi et al. [3] proposed a randomized mechanism that achieves at least half the maximum exchange size and proved that no mechanism can exceed 7/8 of the optimal size under these constraints. Carvalho et al. [21] studied the two-hospital case and characterized Nash equilibria, while Carvalho and Lodi [22] extended these results to social welfare equilibria for an arbitrary number of players, also assuming a bound of 2 on the exchange length, as in our model. These equilibria are compatible with maximum-size solutions, but the authors do not address coalition deviations, which are central to our setting. Gourvès et al. [27] showed that deciding individual rationality in weighted bipartite matchings is NP-hard, illustrating the computational complexity of incentive-related questions in this domain.
Mechanism design for clearinghouses. Roth et al. [40] introduced a priority-based mechanism for pairwise exchanges that is strategy-proof and ensures that revealing all available pairs is a dominant strategy. Sönmez and Ünver [46] generalized this approach, showing that there exist mechanisms that are simultaneously strategy-proof and Pareto-optimal only for pairwise exchanges. However, they also proved that no such mechanism exists once longer cycles or chains are allowed, which fundamentally limits this line of work. Blom et al. [18] proposed a rejection-proof mechanism, meaning that it is a weakly dominant strategy for participants to accept its outcome given that others also accept.
Our contribution and relation to the literature. Our work combines ideas from cooperative and non-cooperative perspectives. We model IKEPs as cooperative games with non-transferable utility, where players (countries or hospitals) seek to maximize the number of their patients who receive transplants. We focus on two stability notions: the strong core and the weak core. Unlike many non-cooperative mechanisms, these notions do not require a loss in efficiency: we show that whenever the strong or weak core is nonempty, there always exists a core solution that achieves the maximum possible number of transplants. Thus, our framework provides a path to designing systems that incentivize participation while fully preserving efficiency. This distinguishes our approach from much of the prior literature, where incentive guarantees often come at the expense of the total number of transplants.
We note that kidney exchange programs can also include altruistic (nondirected) donors – i.e. donors without a related patient in need of a kidney – who can start a chain of transplantations [39, 4]. While our model does not deal with altruistic donors, this is in line with most game-theoretical literature on kidney exchange, both in Europe [8, 9, 10, 31] and in the US [3, 5, 41].
Empirical studies.
Biró et al. [14] provided a comprehensive overview of European kidney exchange programs. Many countries, such as Norway, Denmark, Sweden, and France, only allow pairwise exchanges, while others like the UK permit up to 3-way exchanges. Druzsin et al. [25] conducted simulation studies demonstrating that allowing full cooperation – rather than restricting international exchanges to hard-to-match patients – could lead to significantly more transplants.
Mechanisms based on multiround credit systems have also been studied [28, 31, 34]. In these approaches, cooperative game theory concepts are used to determine target allocations, and matchings in each round are selected to approximate these targets. Benedek et al. [10] conducted large-scale simulations, while [9] analyzed the theoretical complexity of these problems.
1.5 Preliminaries